OpenStudy (anonymous):

Find the length of an arc of 40° in a circle with an 8 inch radius.

OpenStudy (anonymous):

40 degrees is to 360 degrees as L (arc length) is to the circumference of the circle [ 2 pi times 8] 40/ 360 = L / 16 pi 1/9 = L / 50.265 9L = 50.265 L = 5.59 inches approximately

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

ya i tried that formula but the anwser is not on my worksheet

OpenStudy (anonymous):

really? i mean should be it , maybe your teacher wrote something wrong.

OpenStudy (anonymous):

the anwsers are these: 8pi/9 16pi/9 64pi/9 would it be the second i think it is but not sure

OpenStudy (anonymous):

yeay i not sure either. sorry :(

OpenStudy (anonymous):

its alright thx for your ur help tho :)